# Risk-Aware Real-Time Task Allocation for

### v1

@@ Abstract
This paper addresses the control synthesis of heterogeneous stochastic linear multi-agent systems with real-time allocation of signal temporal logic (STL) specifications. Based on previous work, we decompose specifications into sub-specifications on the individual agent level. To leverage the efficiency of task allocation, a heuristic filter evaluates potential task allocation based on STL robustness. Subsequently, an auctioning algorithm determines the definite allocation of specifications. Finally, a control strategy is synthesized for each agent-specification pair using tube-based Model Predictive Control (MPC), ensuring provable probabilistic satisfaction. We demonstrate the efficacy of the proposed methods using a multi-bus scenario that highlights a promising extension to autonomous driving applications like crossing an intersection.
@@ I Introduction
Multi-agent systems are frequently employed in complex tasks that require close collaborations. Formal specifications, such as signal temporal logic (STL), are increasingly utilized to define complex collaborative task objectives [1]. The collaborative control of multi-agent systems satisfying STL specifications has been studied using mixed-integer programming [2], funnel functions [3], and control barrier functions [4]. Nevertheless, a critical challenge for the practical implementation of these methods is their reliability in uncertain and dynamic environments. For systems subject to stochastic disturbances, it is important to evaluate the impact of these disturbances on the task accomplishment of individual agents, so that the associated risks are restricted [5]. When tasks are assigned at runtime [6], task allocation and control become challenging in the multi-agent framework.
@@
Despite recent efforts, there exists a notable gap in the literature regarding the control of stochastic multi-agent systems under real-time allocated STL specifications with probabilistic guarantees. Such systems have significant practical value, exemplified by scenarios such as the two-vehicle intersection case in Figure 1. In such a scenario, each vehicle must navigate while considering its stochastic disturbances and avoiding collisions with its opponent. The complexity of the problem necessitates decomposing the overarching driving specification into two sub-specifications tailored to each vehicle. These sub-specifications, characterized by different time intervals, may require real-time assignment. Hence, this paper addresses the control synthesis of stochastic multi-agent systems under collaborative STL specifications assigned at runtime with probabilistic guarantees.
Figure 1: Illustration of decomposing tasks for autonomous driving.
@@
A crucial aspect of addressing this challenging problem involves task assignment or allocation (TA), which aims to determine optimal agent-task pairs in a multi-agent system [7]. This process yields a linear programming problem  [8], efficiently solvable using auction algorithms [9]. Dynamically assigned tasks have been described using real-time STL specifications [10, 11]. TA necessitates decomposing specifications assigned to multiple collaborative agents into individual agent specifications, requiring the decomposition of STL specifications both among agents [12, 13] and across time [14]. With decomposed real-time specifications, control synthesis for stochastic multi-agent systems holds promise with the utilization of shrinking-horizon model predictive control (MPC) [15] to ensure bounded risks [16]. In previous work, probabilistic reachable tubes have been utilized to optimize risk bounds for both predefined specifications [17] and real-time specifications [6]. These methods underpin the risk-aware control synthesis of stochastic multi-agent systems with decomposed real-time specifications.
@@
This paper proposes a novel approach to allocating STL specifications, involving subgroups of agents, and synthesizing control strategies for individual agents in real-time. By decomposing newly assigned specifications into agent-level tasks, we employ a heuristic approach based on STL robustness [18] to assess agent-specification pairs and attack task assignment via an auctioning scheme formulated as a mixed integer linear program. Subsequently, a tube-based model predictive control strategy is employed where an optimization is solved at every time step at agent level. The solution of this optimization ensures the satisfaction of running and newly allocated tasks with a desired risk level, while discarding tasks not adhering to the risk specification. We prove recursive feasibility of this MPC scheme, leading to an open-loop system with provable probabilistic guarantees and a closed-loop system for which guarantees are suggested.
@@ II Preliminaries and Problem Statement
This section introduces the preliminaries, class of systems, temporal logics, and basic assumptions, culminating in a problem statement and an approach.
II-A Preliminaries
For a given probability measure $\mathbb{P}$ defined over Borel measurable space $(\mathbb{X},\mathcal{B}(\mathbb{X}))$, we denote the probability of an event $\mathcal{A}\in\mathcal{B}(\mathbb{X})$ as $\mathbb{P}(\mathcal{A})$. In this paper, we will work with Euclidean spaces and Borel measurability. Further details on measurability are omitted, and we refer the interested reader to [19]. Additionally, any half-space $H\subset\mathbb{R}^{n}$ is defined as $H:=\{x\in\mathbb{R}^{n}\mid g^{T}x\geq b\}$ with $g\in\mathbb{R}^{n}$ and $b\in\mathbb{R}$. A polyhedron is the intersection of finitely many half-spaces, also denoted as $H:=\{x\in\mathbb{R}^{n}\mid Gx\geq b\}$ with $G\in\mathbb{R}^{q\times n}$ and $b\in\mathbb{R}^{q}$. A polytope is a bounded polyhedron. The 2-norm is defined by $||x||=\sqrt{x^{T}x}$ with $x\in\mathbb{R}^{n}$. The identity matrix is denoted by $I_{n}\in\mathbb{R}^{n\times n}$. The vector of all elements one is denoted by $\bm{1}_{n}\in\mathbb{R}^{n}$, but if the context is clear, it will instead be denoted by $\bm{1}$. The zero vector is denoted by $\bm{0}_{n}\in\mathbb{R}^{n}$, but if the context is clear, it will instead be denoted by $\bm{0}$. For any finite set $I$, the cardinality is denoted by $|I|$. $\mathrm{Proj}_{X}(Y)$ denotes the projection of elements in $Y$ onto $X$.
II-B Multi-agent Systems
@@
For each agent $i\in\mathcal{I}$, we assume there is a local controller. We define the local controllers as a sequence of policies
@@
II-C Signal Temporal Logic & Probabilistic Satisfaction
To mathematically describe tasks, e.g., reaching a target within a given time frame, we consider specifications given by signal temporal logic. Here, we assume that all STL specifications adhere to the negation normal form (NNF), see [1]. This assumption does not restrict the overall framework since [21] proves that every STL specification can be rewritten into the negation normal form.
@@ II-C Signal Temporal Logic & Probabilistic Satisfaction
Here $h_{i}:\mathbb{R}^{nl}\to\mathbb{R}$ indicates the $i$-th element of function $h$. We will assume that all predicate functions are affine functions of the form $h(x)=Gx-b$ with $G\in\mathbb{R}^{q\times nl}$, $b\in\mathbb{R}^{q}$ and $||g_{j}||=1$ for all $j\in\{1,\cdots,q\}$, where $g_{j}$ is the $j$-th row of $G$. Furthermore, we assume that $H_{\mu^{l}}:=\{x\in\mathbb{R}^{nl}\mid Gx\geq b\}$ describes a polytope. Let the set of all predicates be given by $\mathcal{P}$. The STL syntax will be given by
@@
Here, we introduced a slight abuse of the notation as $[a,b]$ is used to describe the set of all natural numbers within the real interval $[a,b]$. Additional operators can be derived such as the eventually-operator $\lozenge_{[a,b]}\phi:=\top U_{[a,b]}\phi$.
@@
For each specification $\phi$, we assume that given is the maximal allowable risk $r_{\phi,\max}$ of not satisfying specification $\phi$. In the remainder, we will refer to this as the maximal risk. Note that for $|\mathcal{I}|=1$, the above definitions of STL and probability will become the standard ones as seen in [6].
II-D Problem Statement & Approach
We first consider an example (see Figure 2) to illustrate the problem and approach covered within this paper. Next, we give a problem description and a problem statement. Lastly, we elaborate upon the approach as seen in Figure 3.
@@ II-D Problem Statement & Approach
Problem Description. Assume each agent $i\in\mathcal{I}$ has dynamics (1) with noise distribution $\mathcal{Q}^{i}_{w}$ and state update $x^{i}(k)$, at each time step $k\in\mathbb{N}$. The objective is to update the suffix controller $\bm{f}^{\mathcal{I}}_{k}$ at each time step $k\in\mathbb{N}$. Here, any newly provided specification is either accepted or rejected based on the maximal risks of the newly provided specification and previously accepted specifications.
@@
Problem Statement. Given a sequence of STL specifications $\bm{\phi}:=\{\phi_{1},\cdots,\phi_{t}\}$ assigned, respectively, at time instances $\{k_{1},\cdots,k_{t}\}$ with $k_{i}<k_{j}$ for $i<j$, with maximal risk $r_{\phi_{i},\max}$, develop a method for updating, at each time step $k\in\mathbb{N}$, suffix $\bm{f}^{\mathcal{I}}_{k}$ such that $\mathbb{P}_{\bm{f}}(\phi_{i},k_{i})\geq 1-r_{\phi_{i},\max}$ if $k_{i}\leq k$, where $\phi_{i}$ with $k_{i}=k$ is allowed to be rejected.
@@
Approach. First, a newly provided STL specification and maximal risk are decomposed onto the single-agent level (Section IV-A). Subsequently, a filter determines which agent-specification pairs are favorable based on STL robustness and heuristics (Section IV-B). Afterwards, for each remaining agent-specification pair, a potential control strategy is determined together with the local risk value (Section III). Based on the local risk value for each feasible agent-specification pair, auctioning determines the definitive assignment (Section IV-C). Should no valid assignment exist, the specification will be rejected. Finally, the control strategy for each agent is updated.
@@
Figure 3: Illustration of the developed approach at time $k$. (0) Determine for each agent-specification pair a control strategy and the local risk value. (1) Auctioning offers from agent-specification pairs. (2) Decision after auctioning. Red is covered in Section III and orange in Section IV.
@@
III Risk-Aware Single-Agent Control Synthesis
Before the multi-agent case, this section considers the single-agent case. Consequently, every specification is on the single-agent level, and only a single agent-specification pair exists. Thus, decomposition of the specification, filtering of the pairs, and auctioning are omitted. Instead, this section will cover control synthesis and local risk value computation for any individual agent-specification pair.
The section will be structured as follows. First, we will consider decomposing the single-agent dynamics, define probabilistic reachable tubes and relate each specification to its maximal risk via (mostly) linear constraints. Second, we utilize the previously mentioned to formulate a tube-based MPC problem to synthesize a control strategy and determine the local risk value for any given agent-specification pair. This section may be regarded as a self-contained summary of our previous work on risk-aware MPC for single-agent systems [6]. We refer the reader to [6] for more details.
III-A Decomposing the Single-Agent Dynamics
In this paper, we decompose dynamics (1), given by
@@ III-A Decomposing the Single-Agent Dynamics
Here $A^{i}_{K}=A^{i}+B^{i}K^{i}$ and $K^{i}$ is a stabilizing feedback gain meant to keep the error $e^{i}$ small. Throughout this section, we omit superscripts indicating agent indices from our notation.
III-B Probabilistic Reachable Tubes
We define probabilistic reachable sets similar to [22].
Definition 2 (Probabilistic Reachable Sets)
A probabili-stic reachable set (PRS) of dynamic probability level $p$ for error dynamics (4b), denoted by $\mathcal{R}^{p}$, is a set that satisfies
$e(0)=\bm{0}\Rightarrow\mathbb{P}(e(k)\in\mathcal{R}^{p})\geq p,\ \forall k\geq 0.$ (5)
Definition 2 states that a PRS of probability level $p$ will contain at any time instance $k$ the accumulating error $e(k)$ with at least probability $p$, if the initial error satisfies $e(0)=\bm{0}$. Multiple representations for probabilistic reachable sets exist, including the ellipsoidal representation. The ellipsoidal representation is obtained from the multivariable Chebyshev inequality; details are given in [22, 23]. For simplicity, we will make the following assumption.
@@ III-B Probabilistic Reachable Tubes
Distribution $\mathcal{Q}_{w}$ has zero mean.
The above assumption is not necessary to obtain an ellipsoidal representation but will simplify most of the computations. Under Assumption 1, the ellipsoidal representation is given by
$\mathcal{R}^{p}=\{x\in\mathbb{R}^{n}\mid x^{T}\Sigma_{\infty}^{-1}x\leq\tilde{p}\},$ (6)
where $\Sigma_{\infty}$ solves the Lyapunov equation $A_{K}\Sigma_{\infty}A_{K}^{T}+\text{var}(\mathcal{Q}_{w})=\Sigma_{\infty}$, $\tilde{p}=\frac{n}{1-p}$, and $p$ is the probability level. Since the variance of $\mathcal{Q}_{w}$ is assumed to be strictly positive definite and $A_{K}$ is stable, Lyapunov theory states that $\Sigma_{\infty}$ is strictly positive definite. In the sequel, we consider the transformed state and constraints for which $\Sigma_{\infty}=I_{n}$. These can be obtained by a simple coordinate transformation $y=\Sigma_{\infty}^{-\frac{1}{2}}x$. We call the transformed dynamics $y$ the normalized dynamics, and in the remainder, we maintain our notations as if the dynamics are normalized.
By defining a PRS $\mathcal{R}^{p_{k}}$ at each time instance $k\in\{0,\cdots,N\}$, we obtain a finite sequence of PRS called a probabilistic reachable tube (PRT) :=$\mathcal{R}:=\{\mathcal{R}^{p_{0}},\cdots,\mathcal{R}^{p_{N}}\}$. We define the risk level $r_{k}=1-p_{k}$ as an upper bound on the probability that $\mathcal{R}^{p_{k}}$ does not contain the error $e(k)$.
III-C Relating Specification and Maximal Risk
In this subsection, we aim to relate each STL specification $\phi$ to its maximal risk $r_{\phi,\max}$ via (mostly) linear constraints. To achieve this, we relate each STL specification $\phi$ assigned at time $k$, by means of a PRT, to the risk $r_{\phi}^{k}$ of failing specification $\phi$ assigned at time $k$. Subsequently, the desired relation can be derived by imposing that $r_{\phi}^{k}\leq r_{\phi,\max}$.
We first define the active horizon $\mathcal{H}_{\phi}^{k}$ of an STL formula $\phi$ as the set of all time instances needed to evaluate $\bm{x}_{k}\vDash\phi$ (see [6, Definition 4]). Here, it should be noted that any active horizon consists of only finite elements. Notice that the suffix $\bm{x}_{N+1}$ with $N=\max{\mathcal{H}_{\phi}^{k}}$, does not influence the satisfaction of $\bm{x}_{k}\vDash\phi$. Hence, in the remainder of this section, we consider only the prefix $\bm{x}_{[k,N]}:=\{x(k),\cdots,x(N)\}$ and note that $\mathbb{P}_{\bm{f},k}({\bm{x}}_{k}\vDash\phi\mid x(k))=\mathbb{P}_{\bm{f},k}({\bm{x}}_{[k,N]}\vDash\phi\mid x(k))$.
Given an STL specification $\phi$ assigned at time $k$, [1] explains that mixed integer linear constraints can represent any STL specification in negation normal form. As can be observed in the following equations, for each time instance $\tau\in\mathcal{H}_{\phi}^{k}$, these STL specifications define a linear constraint that is parametrized with binary variable $s$, as follows
$G(\tau)x(\tau)+MF(\tau)s$ $\geq b(M,\tau)+\rho(\tau)\bm{1},$ (7a)
$Cs\geq d,\ \epsilon\leq\rho(\tau)$ $\leq\frac{M}{2},\ \forall\tau\in\mathcal{H}_{\phi}^{k},$ (7b)
where $G(\tau)$, $E(\tau)$, $C$ and $b(M,\tau)$, $d$ are, respectively, matrices and vectors of appropriate sizes; $M$ is a large positive number; $\epsilon$ is a small positive number; and $\rho(\tau)>0\in\mathbb{R}$ is a positive real number.
We consider the following proposition to quantify risk $r_{\phi}^{k}$. The proof follows from [6, Theorem 5], which utilizes a PRT with risk levels $r(\tau)=\frac{n}{\rho(\tau)^{2}}$.
Proposition 3
Let specification $\phi$ be assigned at time $k$ to the system (4) with normalized dynamics. If the nominal trajectory $\{z(k),\cdots,z(N)\}$ with $z(k)=x(k)$ and nominal controls $\{v(k),\cdots,v(N-1)\}$ satisfies (7) with $\{\rho(k),\cdots,\rho(N)\}$ strictly positive, then
$\mathbb{P}_{\bm{f},k}({\bm{x}}_{k}\vDash\phi\mid x(k))\geq 1-r_{\phi}^{k},$
where risk $r_{\phi}^{k}$ is at most
$r_{\phi}^{k}=\sum_{\tau\in\mathcal{H}_{\phi}^{k}\backslash\{k\}}\frac{n}{\rho(\tau)^{2}}.$ (8)
A relation between specification $\phi$ assigned at time $k$ and risk $r_{\phi}^{k}$ follows from constraints (7) and (8).

### v2

@@ Abstract
This paper addresses the control synthesis of heterogeneous stochastic linear multi-agent systems with real-time allocation of signal temporal logic (STL) specifications. Based on previous work, we decompose specifications into sub-specifications on the individual agent level. To leverage the efficiency of task allocation, a heuristic filter evaluates potential task allocation based on STL robustness, and subsequently, an auctioning algorithm determines the definitive allocation of specifications. Finally, a control strategy is synthesized for each agent-specification pair using tube-based model predictive control (MPC), ensuring provable probabilistic satisfaction. We demonstrate the efficacy of the proposed methods using a multi-shuttle scenario that highlights a promising extension to automated driving applications like vehicle routing.
@@ I Introduction
Multi-agent systems are frequently employed in complex tasks that require close collaborations. Formal specifications, such as signal temporal logic (STL), are increasingly utilized to define complex collaborative task objectives [1]. The collaborative control of multi-agent systems satisfying STL specifications has been studied using mixed-integer programming [2], funnel functions [3], and control barrier functions [4]. Nevertheless, a critical challenge for the practical implementation of these methods is their reliability in uncertain and dynamic environments. For systems subject to stochastic disturbances, it is important to evaluate the impact of these disturbances on the task accomplishment of individual agents so that the associated risks are restricted [5].
@@
Despite recent efforts, a notable gap exists in the literature regarding the control of stochastic multi-agent systems under real-time allocated STL specifications with probabilistic guarantees. Such systems have significant practical value, exemplified by scenarios such as vehicle routing [6]. For example, consider a scenario where two passengers require taxi rides, as illustrated in Fig. 1. This naturally renders a task allocation problem by assigning each passenger a taxi. Finding the most efficient assignment is a critical challenge.
Refer to captionRefer to captionRefer to captionDecompose &AllocateOR
Figure 1: Decomposing and allocating tasks for vehicle routing.
@@
Inspired by the example, this paper addresses 1) the problem of task allocation (TA) of STL specifications, which aims to determine an efficient agent-specification mapping in a multi-agent setting and 2) the control synthesis of agent-specification pairs assigned at runtime with probabilistic guarantees. TA necessitates decomposing specifications assigned to multiple collaborative agents into individual agent specifications, requiring the decomposition of STL specifications both among agents [7, 8] and across time [9]. Previous work yields a linear programming problem [10], efficiently solvable using auction algorithms [11]. With decomposed real-time specifications, control synthesis for stochastic multi-agent systems holds promise with the utilization of shrinking-horizon model predictive control (MPC) [12] to ensure bounded risks [13]. In previous work, probabilistic reachable tubes have been utilized to optimize risk bounds for both predefined specifications [14] and real-time specifications [15]. These methods underpin the risk-aware control synthesis of stochastic multi-agent systems with decomposed real-time specifications.
@@
This paper proposes a novel approach to allocating STL specifications, involving subgroups of heterogeneous agents and synthesizing control strategies for individual agents in real-time. By decomposing newly assigned specifications into agent-level tasks, we employ a heuristic approach based on STL robustness [16] to assess agent-specification pairs and attack task assignment via an auctioning scheme formulated as a mixed integer linear program. Subsequently, a tube-based model predictive control strategy is employed where an optimization is solved at every time step at agent level. The solution of this optimization ensures the satisfaction of previous and newly allocated tasks with a desired risk level, while discarding tasks not adhering to the risk specification. We prove recursive feasibility of this MPC scheme, leading to an open-loop system with provable probabilistic guarantees and a closed-loop system for which probabilistic guarantees follow naturally. As trust is crucial in dynamic multi-agent systems, where agents and environment may often change [17], our contribution is proposing a novel concept of a trustworthy (correct-by-design) control algorithm for multi-agent systems with real-time task assignment.
@@ II Preliminaries and Problem Statement
For a given probability measure $\mathbb{P}$ defined over a Borel measurable space $(\mathbb{X},\mathcal{B}(\mathbb{X}))$, we denote the probability of an event $\mathcal{A}\in\mathcal{B}(\mathbb{X})$ as $\mathbb{P}(\mathcal{A})$. In this paper, we will work with Euclidean spaces and Borel measurability. Further details on measurability are omitted, and we refer the interested reader to [18]. Additionally, any half-space $H\subset\mathbb{R}^{n}$ is defined as $H:=\{x\in\mathbb{R}^{n}\mid g^{T}x\geq b\}$ with $g\in\mathbb{R}^{n}$ and $b\in\mathbb{R}$. A polyhedron is the intersection of finitely many half-spaces, also denoted as $H:=\{x\in\mathbb{R}^{n}\mid Gx\geq b\}$ with $G\in\mathbb{R}^{q\times n}$ and $b\in\mathbb{R}^{q}$. A polytope is a bounded polyhedron. The 2-norm is defined by $||x||=\sqrt{x^{T}x}$ with $x\in\mathbb{R}^{n}$. The identity matrix is denoted by $I_{n}\in\mathbb{R}^{n\times n}$. The vector of all elements one and zero are denoted by $\bm{1}_{n}\in\mathbb{R}^{n}$ and $\bm{0}_{n}\in\mathbb{R}^{n}$, re-spectively. If the context is clear, they will be denoted by $\bm{1}$ and $\bm{0}$. For any finite set $I$, the cardinality is denoted by $|I|$. $\mathrm{Proj}_{X}(Y)$ denotes the projection of elements in $Y$ onto $X$.
II-A Multi-agent Systems
@@
For each agent $i\in\mathcal{I}$, we require there is a local controller. We define the local controllers as a sequence of policies
@@
II-B Signal Temporal Logic & Probabilistic Satisfaction
To mathematically describe tasks, e.g., reaching a target within a given time frame, we consider specifications given by signal temporal logic (STL). Here, we assume that all STL specifications adhere to the negation normal form (NNF), see [1]. This assumption does not restrict the overall framework since [20] shows that every STL specification can be rewritten into the negation normal form.
@@ II-B Signal Temporal Logic & Probabilistic Satisfaction
Here $h_{i}:\mathbb{R}^{nl}\to\mathbb{R}$ indicates the $i$-th element of function $h$. We will assume that all predicate functions are affine functions of the form $h(x)=Gx-b$ with $G\in\mathbb{R}^{q\times nl}$, $b\in\mathbb{R}^{q}$ and $||g_{j}||=1$ for all $j\in\{1,\cdots,q\}$, where $g_{j}$ is the $j$-th row of $G$. Furthermore, we require that $H_{\mu^{l}}:=\{x\in\mathbb{R}^{nl}\mid Gx\geq b\}$ describes a polytope. Let the set of all predicates be given by $\mathcal{P}$. The STL syntax will be given by
@@
Here, we introduced a slight abuse of the notation as $[a,b]$ is used to describe the set of all natural numbers within the real interval $[a,b]$. Note that the definition is invariant of the actual indices of the $l$ agents. Additional operators can be derived such as the eventually-operator $\lozenge_{[a,b]}\phi:=\top U_{[a,b]}\phi$.
@@
For each specification $\phi$, we require that the maximal allowable risk $r_{\phi,\max}$ of not satisfying specification $\phi$ is given. In the remainder, we will refer to this as the maximal risk. Note that for $|\mathcal{I}|=1$, the above definitions of STL and probability will become the standard ones as seen in [15].
II-C Problem Statement & Approach
@@
Consider each agent $i\in\mathcal{I}$ has dynamics (1) with noise distribution $\mathcal{Q}^{i}_{w}$ and state update $x^{i}(k)$, at each time step $k\in\mathbb{N}$. We assume that a new specification may be provided at any time. The objective is to update the suffix controller $\bm{f}^{\mathcal{I}}_{k}$ at each time step $k\in\mathbb{N}$. Here, any newly provided specification is either accepted or rejected based on the maximal risks of the newly provided specification and previously accepted specifications. An illustrative example can be found in the Appendix to further showcase the paper’s objective. We consider the following problem statement.
@@
Problem Statement. Given a sequence of STL specifications $\bm{\phi}:=\{\phi_{1},\cdots,\phi_{t}\}$ assigned, respectively, at time instances $\{k_{1},\cdots,k_{t}\}$ with $k_{i}<k_{j}$ for $i<j$, with maximal risk $r_{\phi_{i},\max}$, develop a method for updating, at each time step $k\in\mathbb{N}$, suffix $\bm{f}^{\mathcal{I}}_{k}$ such that $\mathbb{P}_{\bm{f}}(\phi_{i},k_{i})\geq 1-r_{\phi_{i},\max}$ if $k_{i}\leq k$ and $\phi_{i}$ is a previously accepted specification, where $\phi_{i}$ with $k_{i}=k$ is allowed to be rejected.
@@
Approach. First, a newly provided STL specification and maximal risk are decomposed onto the single-agent level (Section III-B). Subsequently, a filter determines which agent-specification pairs are favorable based on STL robustness and heuristics (Section III-C). Afterwards, for each remaining agent-specification pair, a potential control strategy is determined together with the local risk value (Section III-D). Based on the local risk value for each feasible agent-specification pair, auctioning determines the definitive assignment (Section III-E). Should no valid assignment exist, the specification is rejected. Finally, the control strategy for each agent is updated and implemented (Section III-F).
@@
Figure 2: Illustration of the approach at time $k$. (0) Determine for each agent-specification pair a control strategy and the local risk value. (1) Auctioning offers from agent-specification pairs. (2) Decision after auctioning.
@@
III Risk-Aware Multi-Agent Task Allocation and Control Synthesis
To tackle the problem above, we will first consider agent dynamic decomposition, specification decomposition and heuristic agent-specification filtering. Second, we establish control synthesis and local risk value computation for any individual agent-specification pair. Third, we utilize auctioning, based on the local risk values, to obtain a definitive agent-specification assignment. This section concludes with an implementation algorithm and theoretical analysis of the overall method.
III-A Decomposing Agent Dynamics
We decompose the dynamics (1), given by
@@ III-A Decomposing Agent Dynamics
Here $A^{i}_{K}=A^{i}+B^{i}K^{i}$ and $K^{i}$ is a stabilizing feedback gain meant to keep the error $e^{i}$ small. Throughout this section, we omit superscripts indicating agent indices when appropriate.
III-B Specification Decomposition
Let $\phi$ be a newly assigned STL specification with maximal risk $r_{\phi,\max}$, involving a distinct number of agents, $\nu\leq|\mathcal{I}|$. The objective is to decompose the specification $\phi$ into agent-level sub-specifications, $\phi^{j}$, $j\in\mathcal{I}$, and allocate individual maximal risks, $r_{\phi^{j},\max}$, in accordance to $r_{\phi,\max}$. First, we define $\mathrm{dec}(\phi,\nu)=\wedge_{j=1}^{\nu}\phi^{j}$ as the decomposition of $\phi$ into a conjunctive form, with each conjunct representing a sub-specification, $\phi^{j}$, the satisfaction of which depends on a single agent $j$, and $\nu$ being the number of agents needed to satisfy $\phi$. Second, we assign a maximal risk to $\phi^{j}$, following a union-bound argument, which requires $\sum_{j=1}^{\nu}r_{\phi^{j},\max}\geq r_{\phi,\max}$. In this paper, we select a uniform allocation of risks across sub-specifications, obtaining $r_{\phi^{j},\max}=\frac{r_{\phi,\max}}{\nu}$, for all agents $j\in\{1,\ldots,\nu\}$.
The decomposition $\mathrm{dec}(\phi,\nu)$ acts as a transformation on $\phi$ returning a conjunctive formula, the satisfaction of which implies satisfaction of $\phi$. However, the converse is not necessarily true, implying a lack of completeness. There are a few methods [7, 8, 22] for splitting STL specifications into sub-specifications. In this paper, given a specification $\phi$, we convert Boolean and temporal operators to conjunctive forms based on the heuristics in [22] and decompose predicates following a similar approach to [7]. Therein, predicates are defined on general convex sets, whereas negated forms are not studied. Here, we focus on polytopic settings and allow negations. For our setting, we assume the following.
@@ III-B Specification Decomposition
In view of existing decomposition approaches, where sub-specifications are assigned to subteams of agents, Assumption 1 is restrictive. However, in the setting of polytopic predicates, we can produce agent-level sub-specifications by projecting the associated predicates onto the agents’ state space. Before explaining this, we define the following.
Definition 2
The following lemma shows that any (negated) predicate can be decomposed into a conjunction of agent-level sub-predicates, of which satisfaction implies the satisfaction of the original (negated) predicate. The decomposition of the Boolean and temporal operators is described in [22].
Lemma 3
Let $\mu:=(h(x)\geq\bm{0})$ be a predicate, $H:=\{x\mid h(x)\geq\bm{0}\}$ be a polytope, with $x=[x_{1}^{\intercal}\;\ldots\;x_{\nu}^{\intercal}]^{\intercal}$, $x_{i}\in\mathbb{X}^{i}$, $i\in\{1,\ldots,\nu\}$, and $B_{\mathrm{in}}(H)\subseteq H$ and $B_{\mathrm{out}}(H)\supseteq H$ be the associated orthotopes inscribed in and circumscribing $H$, respectively. Then, it holds that (i) $\mathrm{dec}(\mu,\nu)=\wedge_{i=1}^{\nu}\mu_{\mathrm{in}}^{i}$ satisfies $\mu$, with $\mu_{\mathrm{in}}^{i}:=(x_{i}\in\mathrm{Proj}_{\mathbb{X}^{i}}(B_{\mathrm{in}}(H))$, and (ii) $\mathrm{dec}(\neg\mu,\nu)=\wedge_{i=1}^{\nu}\neg\mu_{\mathrm{out}}^{i}$ satisfies $\neg\mu$, with $\mu_{\mathrm{out}}^{i}:=(x_{i}\in\mathrm{Proj}_{\mathbb{X}^{i}}(B_{\mathrm{out}}(H))$.
Proof:
(i) Let $\mu_{\mathrm{in}}:=(x\in B_{\mathrm{in}}(H))$. By construction, it holds that $B_{\mathrm{in}}(H)\equiv\mathrm{Proj}_{\mathbb{X}^{1}}(B_{\mathrm{in}}(H))\times\cdots\times\mathrm{Proj}_{\mathbb{X}^{\nu}}(B_{\mathrm{in}}(H))$, implying that $\mu_{\mathrm{in}}\equiv\wedge_{i=1}^{\nu}\mu_{\mathrm{in}}^{i}$. Since, $B_{\mathrm{in}}(H)\subseteq H$, $\mu_{\mathrm{in}}$ satisfies $\mu$, which leads to $\wedge_{i=1}^{\nu}\mu_{\mathrm{in}}^{i}$ satisfying $\mu$ as needed. (ii) Let $\mu_{\mathrm{out}}:=(x\in B_{\mathrm{out}}(H))$. Similarly to (i), it holds that $\mu_{\mathrm{out}}\equiv\wedge_{i=1}^{\nu}\mu_{\mathrm{out}}^{i}$, since $B_{\mathrm{out}}(H)\equiv\mathrm{Proj}_{\mathbb{X}^{1}}(B_{\mathrm{out}}(H))\times\cdots\times\mathrm{Proj}_{\mathbb{X}^{\nu}}(B_{\mathrm{out}}(H))$, by construction. Then, we have $\neg\mu_{\mathrm{out}}\equiv\lor_{i=1}^{\nu}\neg\mu_{\mathrm{out}}^{i}$. Since $\wedge_{i=1}^{\nu}\neg\mu_{\mathrm{out}}^{i}$ satisfies $\lor_{i=1}^{\nu}\neg\mu_{\mathrm{out}}^{i}$ and $\neg\mu_{\mathrm{out}}$ satisfies $\neg\mu$, we have that $\wedge_{i=1}^{\nu}\neg\mu_{\mathrm{out}}^{i}$ satisfies $\neg\mu$, completing the proof. ∎
An illustrative example of the specification decomposition method can be found in the Appendix.
III-C Heuristic Agent-Specification Filtering
STL is equipped with robustness metrics for assessing the robust satisfaction of a formula [16]. A scalar-valued function $\rho^{\phi}(\bm{x},k)$ of a signal $\bm{x}$ and time $k$ indicates how robustly a signal $\bm{x}$ satisfies a formula $\phi$ at time $k$. The robustness function is defined recursively in [16, Definition 3].
Let $\phi_{k}$ be a newly assigned STL specification at time $k$, and assume that $\mathrm{dec}(\phi_{k},\nu_{k})=\wedge_{j=1}^{\nu_{k}}\phi_{k}^{j}$. In the following, we require that the total number of agents, denoted by $|\mathcal{I}|$, satisfies $|\mathcal{I}|>\nu_{k}$, otherwise, the following filtering procedure is ignored. Although, $\phi_{k}$ is split into $\nu_{k}$ agent-level sub-specifications, the assignment of $\phi_{k}^{j}$ to a specific agent $i\in\mathcal{I}$ is not determined by $\mathrm{dec}(\phi_{k},\nu_{k})$. In fact, each sub-specification, $\phi_{k}^{j}$, may be assigned to any agent $i\in\mathcal{I}$ resulting in $\nu_{k}|\mathcal{I}|$ agent-specification pairs. To reduce the number of potential pairs, we utilize the following heuristic approach: Based on the quantitative semantics of STL [16], we compute the robustness function of $\phi_{k}^{j}$ over the $i$th agent’s trajectory. This metric indicates roughly the distance between the trajectory of agent $i$ and the space of signals that satisfy $\phi_{k}^{j}$ [23], thereby facilitating a preliminary agent-specification assessment. We approximate the $i$th agent’s trajectory by $\bm{z}^{i,k-1}=\{z_{0}^{i,k-1},z_{1}^{i,k-1},\ldots\}$, $i\in\mathcal{I}$, that is, the signal associated with the (deterministic) nominal dynamics of agent $i$, (4a), computed at time $k-1$, which is available at time $k$. Specifically, for each $j\in\{1,\ldots,\nu_{k}\}$, and $i\in\mathcal{I}$, we efficiently calculate $\rho^{\phi_{k}^{j}}(\bm{z}^{i,k-1},k)$, which requires negligible computational cost. Subsequently, for each $\phi_{k}^{j}$, we collect the largest $\nu_{k}$ values of $\rho^{\phi_{k}^{j}}(\bm{z}^{i,k-1},k)$ among the $|\mathcal{I}|$ options, resulting in $(\nu_{k})^{2}$ agent-specification pairs. Hence, should $v_{k}\ll|\mathcal{I}|$, the filtering significantly reduces potential pairs.